OpenStudy (anonymous):

domain of the function square root 6x-30

OpenStudy (anonymous):

oops make that \[6x-30\geq 0\]

OpenStudy (anonymous):

since you cannot take the square root of a negative number you good from there?

OpenStudy (anonymous):

so its 5 and -5

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the solution is an interval, not two numbers

OpenStudy (anonymous):

(-infinity, 5)U(-5,infitnity)

OpenStudy (anonymous):

\[6x-30\geq 0\] add \(30\) get \[6x\geq 30\] divide get \[x\geq 5\]

OpenStudy (anonymous):

so \[[5,\infty)\]

OpenStudy (anonymous):

one interval, not two there is no \(-5\) in the solution here, as long as i was reading it right \[f(x)=\sqrt{6x-30}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

k good